OpenStudy (anonymous):

Hello! How should i solve this ∫e^(-t)/2) *sin⁡(3t)dt I know what is the result but, i dont know how to get there. Thanks

OpenStudy (anonymous):

, is that \[\int e^{-t/2} sin(3t)dt\]

OpenStudy (anonymous):

?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

by parts

OpenStudy (anonymous):

i have tried, but it always repeat cos(3t)/3, sin(3t)/3 and -2*e^(-1/2) or?

OpenStudy (amistre64):

a = b - a a+a = b 2a = b a = b/2

OpenStudy (amistre64):

if you get a repeat; use it ...

OpenStudy (amistre64):

write it up till your first repeat ... and lets see if we can make this more coherent ...

OpenStudy (anonymous):

sin(3t)=u, e^(-t/2)*dt = du -2*e^(-t/2)*sin(3t) + ∫2*e^(-t/2)*cos(3t)/3*dt

OpenStudy (anonymous):

e^(-t/2)*dt = dv

OpenStudy (amistre64):

good, and one more time should get you back around to the sin e stuff

OpenStudy (anonymous):

yes

OpenStudy (amistre64):

\[\int e^{-t/2} sin(3t)dt=e^{-t/2}*3cos(3t)-\int3cos(3t)*\frac{-e^{-t/2}}{2}dt\] \[\int e^{-t/2} sin(3t)dt3=e^{-t/2}cos(3t)+\frac{3}{2}\left(\int e^{-t/2}cos(3t)dt\right)\] \[\int e^{-t/2} sin(3t)dt=...+\frac{3}{2}\left(-3e^{-t/2}sin(3t)-\frac{3}{2}\int e^{-t/2}sin(3t)dt\right)\] i believe thats going up correctly

OpenStudy (anonymous):

\[\sin (3 t)=\text{Complex}\left[e^{3 i t}\right]\] \[e^{-\frac{t}{2}} e^{3 i t}=e^{\left(-\frac{1}{2}+3 i\right) t}\] \[\int e^{\left(-\frac{1}{2}+3 i\right) t} \, dt\] \[\left(-\frac{2}{37}-\frac{12 i}{37}\right) e^{\left(-\frac{1}{2}+3 i\right) t}=\frac{12}{37} e^{-\frac{t}{2}} \sin (3 t)+\frac{1}{37} (-2) e^{-\frac{t}{2}} \cos (3 t)+\] \[ i \left(\frac{1}{37} (-12) e^{-\frac{t}{2}} \cos (3 t)-\frac{2}{37} e^{-\frac{t}{2}} \sin (3 t)\right)\] only taking the complex part =\[ \left(\frac{1}{37} (-12) e^{-\frac{t}{2}} \cos (3 t)-\frac{2}{37} e^{-\frac{t}{2}} \sin (3 t)\right)\]

OpenStudy (amistre64):

\[\int e^{-t/2} sin(3t)dt=...-\frac{9}{4} \int e^{-t/2}sin(3t)dt\] \[\int e^{-t/2} sin(3t)dt + \frac{9}{4} \int e^{-t/2}sin(3t)dt=...\] \[(1+ \frac{9}{4})\int e^{-t/2} sin(3t)dt=...\] \[\int e^{-t/2} sin(3t)dt=\frac{...}{1+\frac{9}{4}}\]

OpenStudy (anonymous):

but will this ever end?

OpenStudy (amistre64):

that is the end.

OpenStudy (amistre64):

get to the point that it repeats itself; and then use that repetition to solve it

OpenStudy (anonymous):

ok, i will try

OpenStudy (anonymous):

thanks

OpenStudy (amistre64):

spose:\[A=\int e^{-t/2}sin(3t)\ dt\] and you get to the point\[A = f(t)-kA\] this is the repetition, so use it. add kA to each side \[A+kA = f(t)-kA+kA\] \[A+kA = f(t)\cancel{-kA+kA}\] \[(1+k)A = f(t)\] \[A = \frac{f(t)}{1+k}\]